OpenStudy (goldps):

Find the values of x and y, x and y are real numbers 5+(x+2y)i=x+3i

OpenStudy (zzr0ck3r):

\(5+xi+2yi=x+3i\) you need the i terms do match and the non i terms to match so x+2y=3 and x=5 so 5+2y=3 gives y=-1, x=5

OpenStudy (zzr0ck3r):

note \(5+xi+2yi=5+(x+2y)i\)